OpenStudy (nincompoop):

\[y^2+2y=\ln |x| +C\] Given that y(1)=-2 Solve the equation in terms of y

OpenStudy (nincompoop):

I'm trying to solve the differential equation \[ (1+y)* \frac{dy}{dx} =\frac{1}{2x}\] where \( y(1)=-2\) I have separated the equation and taken the derivative on either side but hit a roadblock when I attempt to formalize my response in an explicit form.

OpenStudy (dumbcow):

You dont need explicit form to find constant C \[(-2)^2 +2(-2) = \ln(1) + C\] C = 0 Finally to solve for y, i would use completing the square since y is a quadratic \[(y+1)^2 = \ln x +1\] \[y = \pm \sqrt{\ln x + 1} - 1\]

jimthompson5910 (jim_thompson5910):

y^2+2y = ln(|x|) + C y^2+2y+1 = ln(|x|) + C+1 ... add 1 to both sides to complete the square (y+1)^2 = ln(|x|) + C+1 y+1 = +-sqrt( ln(|x|) + C+1 ) y = -1+-sqrt( ln(|x|) + C+1 )

OpenStudy (nincompoop):

ahhhhhh completing the square =_= Should have thought of that Also only negative squareroot would work with the initial condition

OpenStudy (nincompoop):

\[-2 = - \sqrt{\ln (1) + 1} - 1\]

ganeshie8 (ganeshie8):

Yeah, looks the top part of parabola must be discarded

OpenStudy (mathmate):

Yep, y(x)=-1-sqrt(1+log|x|), with y(1)=-2 and substitute into original, (1+y(x))dy/dx=1/(2x)....yay!